# Clothes-Changing Person Re-Identification with Feasibility-Aware Intermediary Matching

### v1

@@ IV-E Visualization
Visualization of Feature Decoupling Results.  Fig. 6 uses t-SNE [49] tool to visualize the sample distribution of the clothes-relevant and clothes-irrelevant feature spaces. From Fig. 6, we can observe that: (1) Compared with clothes-irrelevant features, samples of the same clothing exhibit a more compact clustering in clothes-relevant feature space. Therefore, by utilizing clothes-relevant features in intermediary matching, we can retrieve same-clothes intermediaries of the same identity more easily than using clothes-irrelevant features. This efficiency arises from the reduced presence of outliers within each clothes cluster. (2) In the clothes-relevant feature space, samples of the same identity remain gathered rather than being split up by samples of other identities. This indicates that our loss function $\mathcal{L}_{re}$ for learning clothes-relevant features can preserve the identity-related clothes information while mining fine-grained clothes cues. Therefore, by using our clothes-relevant feature for intermediary matching, we can preclude noisy intermediaries with same clothes but different identities as query/target. In summary, our feature decoupling method adeptly generates clothes-relevant and clothes-irrelevant features suitable for intermediary matching.

### v2

@@ IV-E Visualization
Visualization of Feature Decoupling Results.  Fig. 6 uses t-SNE [41] tool to visualize the sample distribution of the clothes-relevant and clothes-irrelevant feature spaces. From Fig. 6, we can observe that: (1) Compared with clothes-irrelevant features, samples of the same clothing exhibit a more compact clustering in clothes-relevant feature space (e.g., the features of ID3-C2). Therefore, by utilizing clothes-relevant features in intermediary matching, we can retrieve same-clothes intermediaries of the same identity more easily than using clothes-irrelevant features. This efficiency arises from the reduced presence of outliers within each clothes cluster. (2) In the clothes-relevant feature space, samples of the same identity remain gathered rather than being split up by samples of other identities. Also, for samples from different identities with similar clothes (e.g., ID1-C2 and ID2-C2), there is a clear discrepancy between the features of different identities. This indicates that our clothes label and special designed loss function for learning clothes-relevant features can preserve the identity-related information while mining fine-grained clothes cues. Therefore, by using our clothes-relevant feature for intermediary matching, we can ensure matching same-identity intermediaries in priority, and preclude noisy intermediaries with same clothes but different identities as query. In summary, our feature decoupling method adeptly generates clothes-relevant and clothes-irrelevant features suitable for intermediary matching.
@@ V Conclusion
Acknowledgements
This work is partially supported by Natural Science Foundation of China (NSFC): 62306301, 62376259, 62276246, in part by Fundamental Research Funds for the Central Universities, and in part by the National Postdoctoral Program for Innovative Talents under Grant BX20220310.